OpenStudy (anonymous):

Algerbra 1 finals test? Someone please help me study?

OpenStudy (4everaddicted2anime):

What have you learned so far in algebra 1?

OpenStudy (anonymous):

Equations, solving them, adding and subtracting integers, inequalities, word problems (I suck retriceat those) linear function, and absolute value

OpenStudy (4everaddicted2anime):

What kind of equations? Just linear?

OpenStudy (anonymous):

Linear, finding slope, distributing stuff order of operations and yeah

OpenStudy (anonymous):

I'm not sure if I remember most of the terms I told you about

OpenStudy (4everaddicted2anime):

Okay. I'm going to come up with some equations for you to try. I'll help you with anything you can't remember.

OpenStudy (anonymous):

Thank you :D

OpenStudy (4everaddicted2anime):

Your welcome

OpenStudy (anonymous):

:) =)

OpenStudy (4everaddicted2anime):

Here is what I've come up with so far 1. -8-4 2. 10-(-9) 3. 20+(-30) 4. 100+(-15) 5. 4(5-2)+10((1/2)+1) 6. (5+6)/(2*6-1) 7. 21-2(4+5) 8. 16(5)+(-11)+10

OpenStudy (anonymous):

Are you asking me to solve these?

OpenStudy (4everaddicted2anime):

Yeah

OpenStudy (anonymous):

Kay :)

OpenStudy (anonymous):

1.-12 2.-90 3.-600 4.-1500 5. Idk :( 6.idk 7.8+10*21-2=342 8. Idk

OpenStudy (anonymous):

Is this on all the subjects that I've told you about? Or some of it?

OpenStudy (4everaddicted2anime):

Just some of it

OpenStudy (anonymous):

Oh. Mkaye

OpenStudy (4everaddicted2anime):

2. 10-(-9) The parenthesis around the negative nine are just there to separate the two minus symbols. The word form of it would be ten minus negative nine. The two negatives would cancel each other out and make it 10+9 which equals 19.

OpenStudy (anonymous):

So you don't have to distribute if it looks like that

OpenStudy (4everaddicted2anime):

In this case, no

OpenStudy (4everaddicted2anime):

The same thing with the parenthesis goes with numbers 3 and 4 too. So #3 would be 20-30 which equals -10. #4 would be 100-15 which would be 85.

OpenStudy (anonymous):

Oh

OpenStudy (4everaddicted2anime):

#5. 4(5-2)+10((1/2)+1) First you wpuld need to do the stuff in the parenthesis 4(3)+10(1.5) Then you multiply 12+15 Finally you add to get 27

OpenStudy (anonymous):

Where did you get the 15? Wouldn't that be some decimal?

OpenStudy (4everaddicted2anime):

#6. (5+6)/(2*6-1) Okay. To make this easier lwts split up the numerator and the denominator. The numerator is 5+6. The denominator is 2*6-1. Lets work out the numerator first. 5+6=11 Next, lets work out the denominator. 2*6-1 2*6=12 12-1=11 Now lets put our answers for the numerator and denominator back into the fractional form It turns out to be 11/11 when you divide the numerator and denominator you get an answer of 1.

OpenStudy (4everaddicted2anime):

I got 15 from multiplying 10 and 1.5

OpenStudy (anonymous):

Oh okay yeah ur right

OpenStudy (4everaddicted2anime):

#7. 21-2(4+5) First you do yhe stuff in the parenthesis 21-2(9) Then you multiply 21-18 Finally you subtract these to get 3

OpenStudy (anonymous):

Okay :)

OpenStudy (4everaddicted2anime):

#8. 16*5+(-11)+10 First you multiply 80+(-11)+10 Next you subtract 80 and 11 69+10 Finally you add to get 79

OpenStudy (anonymous):

So do you have to multiply 16 and five before distributing the 11? Or no..

OpenStudy (4everaddicted2anime):

Yes to the first part. You do multiply 16 and 5 first. The parenthesis around the -11 are only there to separate the + and the -.

OpenStudy (4everaddicted2anime):

The next four questions are about finding the linear equation 9. (-4,-7)(8,-13) 10. (-1,3) m=3 11. (1,8)(-2,-1) 12. (-4,-1)(2,2)

OpenStudy (4everaddicted2anime):

I have to go to bed. I will check your answers in the morning